Two independent samples were selected from a population with mean µ=210 and standard deviation s=100. The first sample, Group 1, included 49 patients, and the second sample, Group 2, had 64 patients. Now, we like to estimate the standard deviations of the means in Group 1 and Group 2. Which of the following answers are right?

a Both standard deviations of the means in Group 1 and Group 2 are 100.
b The standard deviations of the means in Group 1 and Group 2 are 100 and 12.5, respectively.
c The standard deviations of the means in Group 1 and Group 2 are 14.3 and 12.5, respectively.
d We don’t have enough information to estimate the standard deviations of the means in Group 1 and Group 2.
Answer #1

Standard deviation for group 1 = 100/V49 -100/7 = 14.3

Stadnard deviation for group 2 = 100/\sqrt{64} = 100/8 = 12.5

Option C is correct.